# A Statistical Framework for Understanding Causal Effects that Vary by Treatment Initiation Time in EHR-based Studies

### v1

@@ 5 Decomposing Sources of Variation
While the projection parameter framework and corresponding model selection procedure outlined in Section 4 provide insight into how treatment effects vary across calendar time, they fail to explain why such changes (if any) are observed. In particular, if the candidate MSM selected is non-constant, suggesting that $\chi_{m}(P)$ varies over time, we would like to ascertain the degree to which underlying changes in treatment efficacy—as opposed to distribution shift in effect modifiers—are responsible.
@@ 5.1 A Cross-Trial Contrast to Remove Distribution Shift
To isolate why treatment effects might vary over calendar time, we begin by defining new contrasts by reweighting mean counterfactual outcomes across eligible populations at different points in time:
@@
The estimand $\chi_{j,m}(P)$ represents the difference in mean counterfactual outcomes at time $m$ in a population where trial baseline covariates are standardized to the covariate distribution of the eligible population at time $j$. First, we note that when $j=m$, this contrast reduces to trial-specific ATE defined in Section 3.2, that is, $\chi_{m,m}(P)=\chi_{m}(P)$. Given the definition of the class of estimands $\chi_{j,m}(P)$, one obtains common populations with which to standardize treatment effects across calendar time. Namely, for $m_{1}\neq m_{2}$, differences between $\chi_{j,m_{1}}(P)$ and $\chi_{j,m_{2}}(P)$ cannot be attributable to covariate shift, as the underlying population on which the causal contrast is defined is the same. We exploit this fact in Section 5.3 to introduce a metric for quantifying effect modification.
In order to identify each $\chi_{j,m}(P)$, we introduce an additional assumption similar to those commonly used in transportability analyses (ung2025, dahabreh2019generalizing, dahabreh2020extending).
@@ 5.2 Estimation of Cross-Trial Effects
Eventually, we will need to estimate $\chi_{j,m}(P)$ for all $M^{2}$ pairs of component target trials. As such, fitting a separate model for each $\xi_{j,m}$ is computationally infeasible. Moreover, conditional density estimation is a challenging statistical problem (wasserman2006all). Because $\xi_{j,m}$ is a ratio of conditional densities, however, we can directly apply Bayes’ rule (sugiyama2012density, benz2025, diaz2023) and leverage pooled models, as with $\widetilde{\mu},\widetilde{\pi}$:
@@
where $T$ denotes the trial membership indicator in full the pooled data set, $\mathcal{D}$. Thus, we have turned the task of conditional density estimation into modeling the trial membership indicator conditional on $\bm{L}$ across all eligible subject-trials. Crucially, this is a classification task well suited for machine learning methods, and we can apply any flexible pooled model $\widetilde{\xi}(m\penalty 10000\ |\penalty 10000\ \bm{\ell})$ to estimate $P(T=m\penalty 10000\ |\penalty 10000\ \bm{L}=\ell,E=1)$.
@@ 5.3 Summarizing Changes in Effect Estimates
Ultimately, our interest in $\chi_{j,m}(P)$ is as a stepping stone towards determining the degree to which changes in calendar time-specific treatment effects are explained by changes in effects modifiers compared to something more fundamental in terms of changes in underlying treatment efficacy. We let $\bm{S}\in\mathbb{R}^{M\times M}$ denote the standardization matrix of all $M^{2}$ cross-trial effects, with $S_{jm}=\chi_{j,m}(P)$, and let $\widehat{\bm{S}}$ denote the estimated standardization matrix with entries $\widehat{\chi}_{j,m}$. To quantify sources of variation in cross-trials effects, we define the following variance-like quantities:
@@ 5.5 Hypothesis Testing for $\theta$
In practice, $H_{0}:\theta\in\{0,1\}$ considers testing at both boundary points for $\theta$, somewhat similar to testing variance components in mixed-effects models (stram1994variance) and as such characterizing the asymptotic distribution of $\widehat{\theta}$ under the null would likely yield a challenging ratio involving mixtures of chi-squared distributions. On the other hand, a full nonparametric bootstrap is computationally infeasible in our setting, given that the size of the pooled dataset $\mathcal{D}$ exceeds several million eligible subject-trials for comparisons involving non-surgical patients (Figure S2). Not only does re-estimation of nuisance functions across bootstrapped datasets pose a large computational burden, but so too does prediction of $\widehat{\mu}_{m}(a_{m},\bm{L}_{j})$ and $\widehat{\xi}_{j,m}(\bm{L}_{m})$ given that the number of predictions required is a $M$-fold larger than the size of $\mathcal{D}$.
Instead, we use an asymptotic version of the parametric bootstrap (van2000asymptotic, EfronTibshirani1994) for $\bm{S}$. In particular, under the conditions of Theorem 4, $\sqrt{n}\bigr(\widehat{\bm{S}}-\bm{S})\overset{d}{\to}\mathcal{N}(0,\bm{\Sigma})$ where $\bm{\Sigma}\in\mathbb{R}^{M^{2}\times M^{2}}$ is the covariance matrix of $\bm{S}$, with entries $\bm{\Sigma}(s_{m_{1}j_{1}},s_{m_{2}j_{2}})=\text{Cov}[\dot{\chi}^{*}_{j_{1},m_{1}}(O;P),\dot{\chi}^{*}_{j_{2},m_{2}}(O;P)]$. Critically, this asymptotic normality for $\bm{S}$ holds even when $\theta$ is at a boundary point. Letting $\widehat{\bm{\Sigma}}_{n}$ denote the empirical covariance matrix of cross-trial influence function contributions, we resample $\widehat{\bm{S}}^{(1)},\dots,\widehat{\bm{S}}^{(B)}\sim\mathcal{N}(\widehat{\bm{S}},\widehat{\bm{\Sigma}}_{n})$. Finally, to approximate the sampling distribution of $\widehat{\theta}$, summarize resamples $\widehat{\bm{S}}^{(1)},\dots,\widehat{\bm{S}}^{(B)}$ by their corresponding variance ratios $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$.
Unfortunately, due to the boundary null, bootstrapped replicates $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$ will never be exactly 0 or 1 and thus 95% percentile-based confidence intervals would never reject $H_{0}$. We follow a similar strategy to that of CavaliereNielsenPedersenRahbek2022, and modify our test to allow for a small margin away from the boundary. For some small $\delta>0$, we adopt the test $H_{0}^{(\delta)}:\theta\in[0,\delta]\cup[1-\delta,1]$ versus $H_{1}^{(\delta)}:\theta\in(\delta,1-\delta)$, and reject $H_{0}^{(\delta)}$ if a 95% percentile-based confidence interval is fully within $(\delta,1-\delta)$. One can interpret $\delta$ as the fraction of variability in entries in $\bm{S}$ for which relative changes in causal effects is deemed negligible. The choice of $\delta$ is highly context dependent, as what constitutes clinically meaningful changes is application specific; in the present context of bariatric surgery, we adopt $\delta=0.05$.
@@ 6.1 Guidance Under No Violations of Assumptions 1-4
To illustrate how model selection and hypothesis testing for $\theta$ can be used together to understand reporting of calendar time-varying effects, we consider two scenarios, one where a constant MSM is selected and a second where a non-constant MSM is selected. First, when the selected MSM is a constant, it is reasonable to report a common effect for the study population, $\psi(m;\widehat{\bm{\beta}})=\widehat{\beta}$, regardless of the value of $\theta$. In some settings, $H_{0}^{(0)}:\theta=0$ might be rejected despite a constant MSM having been selected. Such cases typically arise when differences in estimated treatment effects across calendar time are very small and manifest as random deviations centered around a common effect. Altogether, selecting a constant MSM indicates minimal heterogeneity in the underlying effect over calendar time, suggesting that any observed variation is unlikely to be clinically meaningful.
@@ 6.3 Simulations
Extensive simulations, tied closely to the motivating study of bariatric surgery, which validate the utility of the tools outlined in Sections 4 and 5, are provided in Appendix S3. Simulations highlight how MSM model selection varies by sample size and choice of shrinkage parameter $c$, along with the ability of the inferential procedure described in 5.5 to distinguish between effect modification from underlying variation in treatment efficacy.
7 Application of Framework to Bariatric Surgery
Finally, we return to our motivating study of bariatric surgery and apply the proposed methods to examine trends over calendar time from 2005 to 2011. Analyses were conducted across 16 settings, defined by four post-surgical follow-up times (6 months, 1 year, 2 years, and 3 years) and four treatment comparisons (no surgery versus surgery, no surgery versus RYGB or SG separately, and RYGB versus SG). No SG patients met study eligibility criteria until month $m=30$, so analysis for comparisons involving SG was restricted to trials $m\in\{30,\dots 84\}$.
Random forests were used to estimate all pooled nuisance models, separately for each setting, adjusting for age, BMI, sex, race, KP study site, self-reported smoking status, hypertension, dyslipdeimia, type II diabetes, and prescriptions for antilipemic or antihypertensive medications, or insulin. This set of potential confounders directly followed from those used by arterburn2020. Candidate MSMs for projection included constant, linear, and cubic models, and splines with 2 and 3 internal knots. To select among candidate MSMs, we employed the procedure described in Section 4.4 with $c=0.25$. Additional details on precise covariate and outcome definitions, as well as results for other choices of $c$, are provided in the Supplementary Materials.
@@ 7 Application of Framework to Bariatric Surgery
Based on selected MSMs, surgical patients experienced excess 3-year weight change (relative to non-surgical patients) of -27.4% (-30.1%, -24.8%) at $m=1$ and -18.3% (-18.8%, -17.8%) at $m=84$, reflecting a change of 9.1% estimated treatment effect between 2005 and 2011. This temporal shift indicates that the common effect of –19.6% reported in Section 2, obtained by projection onto a constant MSM, obscures an important aspect understanding the efficacy of bariatric surgery.
Similar changes were observed for RYGB patients, with estimated 3-year weight change differences ranging from -28.3% (-31.3%, -25.3%) at $m=1$ to -23.0% (-23.5%, -22.5%) at $m=84$. As reflected in Figure 3, there is little change in the 3-year RYGB effect estimate after $m=20$. Unsurprisingly then, the common effect estimate of -23.1% reported in Section 2 is comprises a reasonable summary of the treatment effect for a much greater portion of the study period.
@@
Table 2: Summary of model selection results and 95% bootstrapped intervals for $\theta$. The minimizing MSM denotes $\widehat{\psi}^{*}$ minimizing $\widehat{L}(\widehat{\psi}$), while the selected MSM denotes the simplest model within $c=0.25$ weighted standard deviations of the minimizer.
It is not overly surprising that underlying treatment efficacy for bariatric surgery relative to no surgery appears to have “declined” over time, given that the composition of bariatric procedures, specifically the proportion of RYGB and SG, has itself changed over time. Nevertheless, that our method is able to discover this well-explained trend offers evidence of its utility. When making comparisons to non-surgical patients, much of the clinical literature aggregates all surgical procedures under a single label “bariatric surgery” (obrien2018microvascular, Harrington2024, benz2024, madenci2024, haneuse2025, fisher2018association). As such, recognizing that both the meaning and the underlying effectiveness of this aggregated treatment label have changed is particularly relevant. Our results also hint at possible improvements in SG technique and improved weight loss outcomes 6-months post surgery, though such improvements are not observed for longer-term outcomes. One possible explanation is that any underlying changes in technique have helped patients reach their weight-loss nadir more quickly, but have not influenced peak weight-loss or rebound effects (courcoulas2013weight, arterburn2020).
@@ 8 Discussion
One key benefit of the sequential target trial approach considered in this work is that the design directly establishes initiator cohorts (i.e., each with a common time zero), through which calendar time-specific treatment effects can be defined. The proposed methods generalize to other study designs, however, provided that the study entry period can be discretized into sufficiently narrow windows through which one can define calendar time-specific effects. For example, head-to-head comparisons of RYGB and SG do not strictly require a sequence of target trials because the start of follow-up is defined by the date of surgery, and a (matched) cohort design may be a reasonable alternative (McTigue2020, benz2025). Under such designs, one might partition the study entry period into small intervals and define contrasts analogous to $\chi_{m}(P)$ and $\chi_{j,m}(P)$, before applying the ideas introduced in this work. The choice of interval length will be application specific, and depend on both how quickly clinical practice can change as well as the frequency of treatment.
@@
We note that there are natural connections between our work and prior work on violations of the consistency assumption in causal inference (cole2009consistency, vanderweele2009consistency, VanderWeele2013). While RYGB and SG can be framed as multiple versions of bariatric surgery, it is perhaps less natural to view each procedure over time as multiple versions of treatment in the usual sense, in part because it is rarely known when treatments differ, and thus which “version” of treatment was received. In a sense, our setup implicitly defines trial-specific versions of treatment, i.e., through the symbols $A_{m}$ and $a_{m}$. However, we choose to view counterfactual outcomes $Y_{m}(a_{m})$ similarly to those defined and framed in cole2009consistency, $Y(a,k)$, where the additional treatment details $(k)$ required to maintain consistency are entirely captured by trial index ($m$) in our notation. In particular, we adopt the viewpoint that if $a_{m}=a_{j}$ but $Y_{m}(a_{m})\neq Y_{j}(a_{j})$, then this reflects changes in the efficacy of a particularly treatment $a$ between two time points $m,j$, rather than always calling this phenomenon multiple versions of treatment. This choice is in line with Cole and Frangakis, who note that taking versions of treatment to the extreme precludes any sort of generalizability and hampers the utility of a study.
Finally, we have assumed throughout that during the time period where a patient is enrolled in Kaiser Permanente insurance, and thus is accessible in the EHR, there is no missing data in components of $O$. In practice, missing covariates may preclude ascertainment of eligibility, or outcomes may be missing even when eligibility is ascertainable. For the purposes of this work, we treat patients with incomplete information at a particular trial as ineligible $(E_{m}=0)$. However, we see combining ideas from our framework with recent work on causal inference under missing eligibility (benz2024, benz2025) as key area of future research. Simultaneously handling possibly missing treatment status and/or outcomes across trials represents another important line of further methodological development.

### v2

@@ 5 Decomposing Sources of Variation
While the projection parameter framework and corresponding MSM selection procedure outlined in Section 4 provide insight into how treatment effects vary across calendar time, they fail to explain why such changes (if any) are observed. In particular, if the candidate MSM selected is non-constant, suggesting that $\chi_{m}(P)$ varies over time, we would like to ascertain the degree to which underlying changes in treatment efficacy—as opposed to distribution shift in effect modifiers—are responsible.
@@ 5.1 A Cross-Trial Contrast to Remove Distribution Shift
To isolate why treatment effects might vary over calendar time, we begin by defining new contrasts by averaging reweighted conditional average treatment effects (CATE) Dahabreh et al. (2019), Cole and Stuart (2010) over covariate distributions reflecting trial eligible populations at different points in time:
@@
The estimand $\chi_{j,m}(P)$ represents the difference in mean counterfactual outcomes among patients eligible for the $m^{\text{th}}$ trial in a population where trial baseline covariates are standardized to the covariate distribution of the eligible population at time $j$. First, we note that when $j=m$, this contrast reduces to trial-specific ATE defined in Section 3.2, that is, $\chi_{m,m}(P)=\chi_{m}(P)$. Given the definition of the class of estimands $\chi_{j,m}(P)$, one obtains common populations with which to standardize treatment effects across calendar time. Namely, for $m_{1}\neq m_{2}$, differences between $\chi_{j,m_{1}}(P)$ and $\chi_{j,m_{2}}(P)$ cannot be attributable to covariate shift, as the underlying population on which the causal contrast is defined is the same. We exploit this fact in Section 5.3 to introduce a metric for quantifying effect modification.
In order to identify each $\chi_{j,m}(P)$, we introduce an additional assumption similar to those commonly used in transportability analyses (Dahabreh et al., 2019, Cole and Stuart, 2010).
@@ 5.2 Estimation of Cross-Trial Effects
Eventually, we will need to estimate $\chi_{j,m}(P)$ for all $M^{2}$ pairs of component target trials. As such, fitting a separate model for each $\xi_{j,m}$ is computationally infeasible. Moreover, conditional density estimation is a challenging statistical problem (Wasserman, 2006). Because $\xi_{j,m}$ is a ratio of conditional densities, however, we can directly apply Bayes’ rule (Sugiyama et al., 2012, Benz et al., 2025b, Díaz et al., 2023) and leverage pooled models, as with $\widetilde{\mu},\widetilde{\pi}$:
@@
where $T$ denotes the trial membership indicator in the full pooled data set, $\mathcal{D}$. Thus, we have turned the task of conditional density estimation into modeling the trial membership indicator conditional on $\bm{L}$ across all eligible subject-trials. Crucially, this is a classification task well suited for machine learning methods, and we can apply any flexible pooled model $\widetilde{\xi}(m\penalty 10000\ |\penalty 10000\ \bm{\ell})$ to estimate $P(T=m\penalty 10000\ |\penalty 10000\ \bm{L}=\ell,E=1)$.
@@ 5.3 Summarizing Changes in Effect Estimates
Ultimately, our interest in $\chi_{j,m}(P)$ is as a stepping stone towards determining the degree to which changes in calendar time-specific treatment effects are explained by changes in effect modifiers compared to something more fundamental in terms of changes in underlying treatment efficacy. We let $\bm{S}\in\mathbb{R}^{M\times M}$ denote the standardization matrix of all $M^{2}$ cross-trial effects, with $S_{jm}=\chi_{j,m}(P)$, and let $\widehat{\bm{S}}$ denote the estimated standardization matrix with entries $\widehat{\chi}_{j,m}$. To quantify sources of variation in cross-trial effects, we define the following variance-like quantities:
@@ 5.5 Hypothesis Testing for $\theta$
In practice, $H_{0}:\theta\in\{0,1\}$ considers testing at both boundary points for $\theta$, somewhat similar to testing variance components in mixed-effects models (Stram and Lee, 1994) and as such characterizing the asymptotic distribution of $\widehat{\theta}$ under the null would likely yield a challenging ratio involving mixtures of chi-squared distributions. On the other hand, a full nonparametric bootstrap is computationally infeasible in our setting, given that the size of the pooled dataset $\mathcal{D}$ exceeds several million eligible subject-trials for comparisons involving non-surgical patients (Figure S2). Not only does re-estimation of nuisance functions across bootstrapped datasets pose a large computational burden, but so too does prediction of $\widehat{\mu}_{m}(a_{m},\bm{L}_{j})$ and $\widehat{\xi}_{j,m}(\bm{L}_{m})$ given that the number of predictions required is a $M$-fold larger than the size of $\mathcal{D}$.
Instead, we use an asymptotic version of the parametric bootstrap (Van der Vaart, 2000) for $\bm{S}$. In particular, under the conditions of Theorem 4, $\sqrt{n}\bigr(\widehat{\bm{S}}-\bm{S})\overset{d}{\to}\mathcal{N}(0,\bm{\Sigma})$ where $\bm{\Sigma}\in\mathbb{R}^{M^{2}\times M^{2}}$ is the covariance matrix of $\bm{S}$, with entries $\bm{\Sigma}(s_{m_{1}j_{1}},s_{m_{2}j_{2}})=\text{Cov}[\dot{\chi}^{*}_{j_{1},m_{1}}(O;P),\dot{\chi}^{*}_{j_{2},m_{2}}(O;P)]$. Critically, this asymptotic normality for $\bm{S}$ holds even when $\theta$ is at a boundary point. Letting $\widehat{\bm{\Sigma}}_{n}$ denote the empirical covariance matrix of cross-trial influence function contributions, we resample $\widehat{\bm{S}}^{(1)},\dots,\widehat{\bm{S}}^{(B)}\sim\mathcal{N}(\widehat{\bm{S}},\widehat{\bm{\Sigma}}_{n})$. Finally, to approximate the sampling distribution of $\widehat{\theta}$, summarize resamples $\widehat{\bm{S}}^{(1)},\dots,\widehat{\bm{S}}^{(B)}$ by their corresponding variance ratios $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$.
Unfortunately, due to the boundary null, bootstrapped replicates $\widehat{\theta}^{(1)},\dots,\widehat{\theta}^{(B)}$ will never be exactly 0 or 1 and thus 95% percentile-based confidence intervals would never reject $H_{0}$. We follow a similar strategy to that of Cavliere et al. Cavaliere et al. (2022), and modify our test to allow for a small margin away from the boundary. For some small $\delta>0$, we adopt the test $H_{0}^{(\delta)}:\theta\in[0,\delta]\cup[1-\delta,1]$ versus $H_{1}^{(\delta)}:\theta\in(\delta,1-\delta)$, and reject $H_{0}^{(\delta)}$ if a 95% percentile-based confidence interval is fully within $(\delta,1-\delta)$. One can interpret $\delta$ as the fraction of variability in entries in $\bm{S}$ for which relative changes in causal effects are deemed negligible. The choice of $\delta$ is highly context dependent, as what constitutes clinically meaningful changes is application specific; in the present context of bariatric surgery, we adopt $\delta=0.05$.
@@ 6.1 Guidance Under No Violations of Assumptions 1-4
To illustrate how MSM selection and hypothesis testing for $\theta$ can be used together to understand reporting of calendar time-varying effects, we consider two scenarios, one where a constant MSM is selected and a second where a non-constant MSM is selected. First, when the selected MSM is a constant, it is reasonable to report a common effect for the study population, $\psi(m;\widehat{\bm{\beta}})=\widehat{\beta}$, regardless of the value of $\theta$. In some settings, $H_{0}^{(0)}:\theta=0$ might be rejected despite a constant MSM having been selected. Such cases typically arise when differences in estimated treatment effects across calendar time are very small and manifest as random deviations centered around a common effect. Altogether, selecting a constant MSM indicates minimal heterogeneity in the underlying effect over calendar time, suggesting that any observed variation is unlikely to be clinically meaningful.
@@ 6.3 Simulations
Extensive simulations, tied closely to the motivating study of bariatric surgery, which validate the utility of the tools outlined in Sections 4 and 5, are provided in Appendix S3. Simulations highlight how MSM selection varies by sample size and choice of shrinkage parameter $c$, along with the ability of the inferential procedure described in Section 5.5 to distinguish between effect modification from underlying variation in treatment efficacy.
7 Evaluating Temporal Trends in Causal Effects of Bariatric Surgery
Finally, we return to our motivating study of bariatric surgery and leverage the proposed methods to examine trends over calendar time from 2005 to 2011. Analyses were conducted across 16 settings, defined by four post-surgical follow-up times (6 months, 1 year, 2 years, and 3 years) and four treatment comparisons (no surgery versus surgery, no surgery versus RYGB or SG separately, and RYGB versus SG). No SG patients met study eligibility criteria until month $m=30$, so analysis for comparisons involving SG was restricted to trials $m\in\{30,\dots,84\}$.
Random forests were used to estimate all pooled nuisance models, separately for each setting, adjusting for age, BMI, sex, race, KP study site, self-reported smoking status, hypertension, dyslipidemia, type II diabetes, and prescriptions for antilipemic or antihypertensive medications, or insulin. This set of potential confounders directly followed from those used by Arterburn et al. Arterburn et al. (2020). Candidate MSMs for projection included constant, linear, and cubic models, and splines with 2 and 3 internal knots. To select among candidate MSMs, we employed the procedure described in Section 4.5 with $c=0.25$. Additional details on precise covariate and outcome definitions, as well as results for other choices of $c$, are provided in the Supplementary Materials.
@@ 7 Evaluating Temporal Trends in Causal Effects of Bariatric Surgery
Based on selected MSMs, surgical patients experienced excess 3-year weight change (relative to non-surgical patients) of -27.4% (-30.1%, -24.8%) at $m=1$ and -18.3% (-18.8%, -17.8%) at $m=84$, reflecting a change of 9.1% in the estimated treatment effect between 2005 and 2011. This temporal shift indicates that the common effect of –19.6% reported in Section 2, obtained by projection onto a constant MSM, obscures an important aspect of understanding the efficacy of bariatric surgery.
Similar changes were observed for RYGB patients, with estimated 3-year weight change differences ranging from -28.3% (-31.3%, -25.3%) at $m=1$ to -23.0% (-23.5%, -22.5%) at $m=84$. As reflected in Figure 3, there is little change in the 3-year RYGB effect estimate after $m=20$. Unsurprisingly then, the common effect estimate of -23.1% reported in Section 2 comprises a reasonable summary of the treatment effect for a much greater portion of the study period.
@@
Table 2: Summary of MSM selection results and 95% bootstrapped intervals for $\theta$. The minimizing MSM denotes $\widehat{\psi}^{*}$ minimizing $\widehat{L}(\widehat{\psi}$), while the selected MSM denotes the simplest model within $c=0.25$ weighted standard deviations of the minimizer.
It is not overly surprising that underlying treatment efficacy for bariatric surgery relative to no surgery appears to have “declined” over time, given that the composition of bariatric procedures, specifically the proportion of RYGB and SG, has itself changed over time. Nevertheless, that our method is able to discover this well-explained trend offers evidence of its utility. When making comparisons to non-surgical patients, much of the clinical literature aggregates all surgical procedures under a single label “bariatric surgery” (O’Brien et al., 2018, Harrington et al., 2024, Benz et al., 2025a, Madenci et al., 2024, Haneuse et al., 2025, Fisher et al., 2018). As such, recognizing that both the meaning and the underlying effectiveness of this aggregated treatment label have changed is particularly relevant. Importantly, while the $m=84$ estimate of -18.3% is more reflective of current real-world efficacy of bariatric surgery than the common effect of -19.6%, differences are attributable to both changing distribution of procedures and patient populations being treated. As such additional standardization efforts may be required to get a truly contemporary estimate, as the DURABLE database only follows patients through 2015.
Finally, results also hint at possible improvements in SG technique and improved weight loss outcomes 6 months post-surgery, though such improvements are not observed for longer-term outcomes. One possible explanation is that any underlying changes in technique have helped patients reach their weight-loss nadir more quickly, but have not influenced peak weight-loss or rebound effects.
@@ 8 Discussion
One key benefit of the sequential target trial approach considered in this work is that the design directly establishes initiator cohorts (i.e., each with a common time zero), through which calendar time-specific treatment effects can be defined. The proposed methods generalize to other study designs, however, provided that the study entry period can be discretized into sufficiently narrow windows through which one can define calendar time-specific effects. For example, head-to-head comparisons of RYGB and SG do not strictly require a sequence of target trials because the start of follow-up is defined by the date of surgery, and a (matched) cohort design may be a reasonable alternative (McTigue et al., 2020, Benz et al., 2025b). Under such designs, one might partition the study entry period into small intervals and define contrasts analogous to $\chi_{m}(P)$ and $\chi_{j,m}(P)$, before applying the ideas introduced in this work. The choice of interval length will be application specific, and depend on both how quickly clinical practice can change as well as the frequency of treatment.
The decomposition framework introduced in Section 5 shares conceptual motivation with Gilbert et al. (Gilbert et al., 2025), who propose a related decomposition to explain heterogeneity in treatment effects across distinct studies or sites. However, the two frameworks differ in important ways. Gilbert et al. decompose between-study heterogeneity into contributions from case-mix differences, mediating variables, and pure effect modification across externally defined, unordered study indices. By contrast our framework is specifically designed for the temporally ordered, calendar time setting, where the goal is to disentangle changes in treatment efficacy from covariate shift occurring within a single EHR-based study. Additionally, rather than estimating decomposition components separately, we introduce a scalar summary metric $\theta$ to quantify the aggregate role of covariate shift in explaining observed temporal variation in treatment effects.
@@
We note that there are natural connections between our work and prior work on violations of the consistency assumption in causal inference (Cole and Frangakis, 2009, VanderWeele, 2009, VanderWeele and Hernán, 2013). While RYGB and SG can be framed as multiple versions of bariatric surgery, it is perhaps less natural to view each procedure over time as multiple versions of treatment in the usual sense, in part because it is rarely known when treatments differ, and thus which “version” of treatment was received. In a sense, our setup implicitly defines trial-specific versions of treatment, i.e., through the symbols $A_{m}$ and $a_{m}$. However, we choose to view counterfactual outcomes $Y_{m}(a_{m})$ similarly to those defined and framed in Cole and Frangakis Cole and Frangakis (2009), $Y(a,k)$, where the additional treatment details $(k)$ required to maintain consistency are entirely captured by trial index ($m$) in our notation. In particular, we adopt the viewpoint that if $a_{m}=a_{j}$ but $Y_{m}(a_{m})\neq Y_{j}(a_{j})$, then this reflects changes in the efficacy of a particular treatment $a$ between two time points $m,j$, rather than always calling this phenomenon multiple versions of treatment. This choice is in line with Cole and Frangakis, who note that taking versions of treatment to the extreme precludes generalizability and hampers the utility of a study. For instance, we would find it counterproductive to define separate versions of bariatric surgery for each operating surgeon, or of non-surgical weight management based on prevailing pharmacological or diet and exercise standards.
Finally, we have assumed throughout that during the time period where a patient is enrolled in Kaiser Permanente insurance, and thus is accessible in the EHR, there is no missing data in components of $O$. In practice, missing covariates may preclude ascertainment of eligibility, or outcomes may be missing even when eligibility is ascertainable. For the purposes of this work, we treat patients with incomplete information at a particular trial as ineligible $(E_{m}=0)$. However, we see combining ideas from our framework with recent work on causal inference under missing eligibility (Benz et al., 2025a, b) as a key area of future research. Simultaneously handling possibly missing treatment status and/or outcomes across trials represents another important line of further methodological development.